OpenStudy (anonymous):

sqrt 300 = 10 sqrt x = 5 sqrt y = 2sqrt k

OpenStudy (anonymous):

can you help me please !

OpenStudy (anonymous):

sqrt300 = sqrt(3*10*10)=10sqrt 3 now you have sqrt 300 = 10 sqrt x so x=3

OpenStudy (anonymous):

sqrt 300 = sqrt(3*2*2*5*5)=5 sqrt 12 but we have sqrt 300 = 5sqrt y hence y=12

OpenStudy (anonymous):

sqrt 300 = sqrt(3*2*2*5*5) = 2 *sqrt 75 but we have sqrt 300 =2 sqrt k hence k=75

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

welcome dear